OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 15 degrees Celsius is.

OpenStudy (anonymous):

multiply through by 9 9C = 5(F - 32) F - 32 = 9C / 5 F = (9C/5) + 32

OpenStudy (anonymous):

part 2 F = (9*15/ 5) + 32 = 27 + 32 = 59

OpenStudy (anonymous):

Thank you ! : )

OpenStudy (anonymous):

yw